OpenStudy (anonymous):

For b(x) = x + 1 and k(x) = 2x – 1, select the true statement. Answer -1/2 is excluded from the domain of (b / k)(x). -1/2 is excluded from the domain of (k / b)(x). -1 is excluded from the domain of (b / k)(x). -1 is excluded from the domain of (k / b)(x).

OpenStudy (anonymous):

helppp

OpenStudy (anonymous):

Do you know function operations? If not: \[(f / g)(x) = (\frac{f}{g})(x)\] See if that helps.

OpenStudy (anonymous):

no i dont know

OpenStudy (anonymous):

Oh crap, sorry, I meant: \[(f / g)(x) = \left(\frac{f(x)}{g(x)}\right)\]

OpenStudy (anonymous):

so x+1/2x+1

OpenStudy (anonymous):

Yes that's \((b / k)(x)\).

OpenStudy (anonymous):

so is the answer B.

OpenStudy (anonymous):

srry C right

OpenStudy (anonymous):

The function in c is \((b / k)(x) \), so \(\frac{x+1}{2x-1}\).

OpenStudy (anonymous):

But d is \((k / b)(x) = \frac{2x-1}{x+1}\), which does exclude \(-1\) in its domain.